Compute the least square regression line for a sample of 6 pairs of observations, given that

xi = 5, 4, 6, 3, 8, 7,    Ʃy = 150,  Ʃxy = 404,  (y-ybar)2=37, (y-yhat)2=15

Also calculate the Co-efficient of determination and interpret the result.


1
Expert's answer
2021-01-14T17:41:07-0500

We obtain the parameters of the least square equation by solving the simultaneous equations


"\sum y_i= m\sum x_i + c"

"\sum x_i y_i= m\sum x_i ^2 + c \sum x_i"


"\sum x_i = 5+4+6+3+8+7=33"

"\sum x_i ^2= 5^2+4^2+6^2+3^2+8^2+7^2=199"


Therefore, we have

"150=33m + c" ------- equation i

"404=199m + 33c" -------- equation ii


From equation i

"c = 150-33m"


Substitute c in equation ii

"404=199m-33(150-33m)"

"404= 199m + 1089m - 4950"

"1288m = 5354"

"m=\frac{5353}{1288}=4.1568"


Therefore

"c=\frac{404-199(4.1568)}{33}=-12.8245"


The least square equation is given as

"y=4.1568x-12.8245"


Coefficient of determination

"R^2 = 1- \frac{RSS}{TSS}=1-\frac{15}{37}= 0.5944"


This means that the random variable x explains 59.44% of the variation in variable y. Other factors not considered in the least square regression are responsible for the remaining 40.56% of the variations in y